What is acceleration?


The change in motion of an object.


The change in speed and/or direction of an object.


The change in speed of an object.


The speed of an object.


1
Expert's answer
2021-03-26T11:40:17-0400

Acceleration is the rate of change of velocity or speed, or a physical quantity that defines how quickly speed or velocity changes:


"a=\frac{\Delta v}{\Delta t}=\frac{v_2-v_1}{t_2-t_1}."

For instance, an RC car increased the speed from 0 (rest) to 20 m/s in just 20 seconds, then the acceleration will be

"a=\frac{20-0}{20}=1\text{ m/s}^2."

As we see, acceleration is measured in meters per second squared.